A car initially traveling at 43 m/s accelerates to 53 m/s. While speeding up, the car travels 200 m. Calculate the acceleration.
1
Expert's answer
2020-10-01T10:10:03-0400

We can find the acceleration of the car from the kinematic equation:


"v_f^2 = v_i^2 + 2as,"

here, "v_i = 43 \ \dfrac{m}{s}" is the initial velocity of the car, "v_f = 53 \ \dfrac{m}{s}" is the final velocity of the car, "s = 200 \ m" is the distance traveled by the car and "a" is the acceleration of the car.

Then, we get:


"a = \dfrac{v_f^2 - v_i^2}{2s} = \dfrac{(53 \ \dfrac{m}{s})^2 - (43 \ \dfrac{m}{s})^2}{2 \cdot 200 \ m} = 2.4 \ \dfrac{m}{s^2}."

Answer:

"a = 2.4 \ \dfrac{m}{s^2}."